OpenStudy (anonymous):

Which expressions are equal to 8 when x = 6 and y = 2? Choose all answers that are correct. A. x – y + 4 B. y + x – 4 C. xy – 2 D. 5 + (x ÷ y)

OpenStudy (y1066380):

Put x=6 and y=2 into the equations and check which of the answers equal to 8

OpenStudy (anonymous):

Which do you think it is? You just need to plug in each value for x and y and see which one equals 8.

OpenStudy (anonymous):

omg ty u guys helped me lots being homeschooled isnt easy for mr its my first yr

OpenStudy (anonymous):

No problem :) Did you get the answer?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Which one do you think it is? (just to see if we got the same answer) :)

OpenStudy (anonymous):

i just dont get c

OpenStudy (anonymous):

a and d

OpenStudy (anonymous):

but i dont know bout c

OpenStudy (y1066380):

How about D?

OpenStudy (anonymous):

You would plug in 6 for x and 2 for y. The equation would be 12-2, which doesn't equal 8, so C is not correct.

OpenStudy (anonymous):

yea i said A and D

OpenStudy (anonymous):

TY!!!!!!!!!!!

OpenStudy (y1066380):

Yes

OpenStudy (anonymous):

A would be 6 + 2 + 4, so that wouldn't equal 8. Are you sure you read it right?

OpenStudy (anonymous):

OH NO I ALREADY ENTERED IT

OpenStudy (y1066380):

For (A), x-y+4 =6-2+4 =8

OpenStudy (y1066380):

A and D is also correct

OpenStudy (anonymous):

Oh, sorry I read that wrong. I read plus not minus. you were right!